ou harve 1 day remaining in your WebAasign Orace Parlod Get Access Todey What is the herizental distance x between the two balloons? 48.2m Section 1.4 MacBook Air 3 5 9 G H
Notes C CJ10 1.P.018 Hint: Consider using the law of cosines.) What is the value of each of the angles of a triangle whose sides are a -90, b- 145, and c - 172 cm in length? tional Materials Section 1.4
Answer #1

balloon problem

tan (theta) = ( h - 48.2) / x

tan 14.7 = (61.5 - 48.2)/x

x = 50.7 m

==========

as per law of cosine

cos (alpha) = ( b^2 + c^2 - a^2)/ ( 2 b*c)

cos alpha = ( 145^2 + 172^2 - 90^2) /( 2 * 145* 172)

alpha = 31.545

Similarly

cos (beta) = ( a^2 + c^2 - b^2)/( 2 a c)

cos (beta) = ( 90^2 + 172^2 - 145^2) / (2* 90*172)

beta = 57.44

again

cos( gamma) = ( a^2 + b^2 - c^2) / ( 2 ab)

cos ( gamma) = ( 90^2 + 145^2 - 172^2) / ( 2* 90* 145)

gamma = 91.015

========

Do comment in case any doubt, will reply for sure.. Goodluck